OpenStudy (anonymous):

Solve each system of equations by substitution -3v-3x=15 5v+x=-29

OpenStudy (tkhunny):

Divide the first by 3 and something magic may pop out.

OpenStudy (skullpatrol):

I tried to teach you this yesterday and you refused to learn?

OpenStudy (anonymous):

I did not refuse I just didn't get it

OpenStudy (tkhunny):

Well, @overtonka, I think @skullpatrol would be delighted to see you make a better effort, today. Divide the first equation by 3 and show what you get.

OpenStudy (anonymous):

-1v-1x=5

OpenStudy (tkhunny):

Super! Any suggestion what we should do next? Hint: I didn't have you divide by 3 just for fun.

OpenStudy (anonymous):

solve for one of the variables?

OpenStudy (skullpatrol):

Yes!

OpenStudy (anonymous):

so 1v=5-1x?

OpenStudy (skullpatrol):

Now what (hint: use the name of the method)

OpenStudy (tkhunny):

Since we are using substitition, that would be the things to do. Please abandon those silly-looking '1's. We don't usually write those unless we are trying to make a point.

OpenStudy (anonymous):

ok so v-x=5?

OpenStudy (tkhunny):

No, why did you UNsolve? Please "solve for one of the variables".

OpenStudy (anonymous):

v=5-x?

OpenStudy (skullpatrol):

Correct.

OpenStudy (anonymous):

so plug that in to the other equation?

OpenStudy (skullpatrol):

Substitution principle: An expression maybe replaced by another expression that has the same value If $$a=b$$then b maybe substituted for a in any expression.

OpenStudy (skullpatrol):



OpenStudy (tkhunny):

It's your big chance. Let's see the substitution!